# Situationally-Aware Dynamics Learning

### v1

@@ II-E Unstructured Terrain Navigation
The validation of our approach is through the challenging task of unstructured terrain navigation, where unmodeled and unmeasured terrain characteristics can significantly impact the robot’s motion. Unstructured terrain navigation has been tackled mostly from two perspectives: adaptation and path planning. Works addressing this from the adaptive perspective typically devise methods that allow the robot to recover from unexpected or hard scenarios. For example, Xu et al. [96] proposes a method to bypass the precise modeling of the state transition function of an MDP with continuous state space and shows that it can be applied for off-road navigation and recovering from unexpected phenomena (e.g., wheel slippage). Wang et al. [94] learns a probabilistic dynamics model trained to consider the propagating effects of uncertainty and uses it along a tracking controller and constrained optimization to find trajectories with low divergence. Terrain-aware methods have been recently developed to generate consistent navigation behaviors by learning a control offset from demonstrations [83, 84]. On the other hand, the methods concerned with global planning typically use the robot’s sensors to analyze the terrain and try to find the trajectory that seems safer and easier to navigate [100, 40, 60]. Other methods have done so by focusing on risk-awareness [81, 10, 11]. Learning-based approaches have considered ground-robot interaction and surface data [77]. Our method can be framed at the intersection of both perspectives. In particular, we obtain dynamics models with greater robustness to terrain challenges and use those to make local motion plans that help the robot traverse safely under partial state information.
@@ III-A Generalized Hidden Parameter Markov Decision Process (GHP-MDP)
Markov Decision Processes (MDPs) provide a framework for modeling decision-making with uncertainty over the action outcomes due to an agent’s interaction with its environment [89]. Hidden Parameter MDPs (HiP-MDP) extend MDPs to scenarios where predicting the future also requires a set of unobservable parameters $\Theta$ that correspond to the sufficient statistics to adapt to a new situation [22]. In HiP-MDPs, $\Theta$ explicitly represent a critical variable of the environment; for example, physical properties like the terrain friction. This makes it necessary to know a-priori which missing variables need to be estimated by $\Theta$, which is a strong assumption we cannot fulfill when the robot is navigating unstructured terrains as we cannot anticipate every challenge it will face.
[70] build upon HiP-MDPs and present GHP-MDPs to account for unobserved, latent parameters which influence the environment’s dynamics and reward function. These latent factors are not directly observable by the agent, but the agent must infer them through interaction with the environment to optimize its policy.
@@ III-A1 Comparison to other MDP Models:
MDP [7]: The state $s$ is fully observable, and the environment’s dynamics are completely determined by the state and action, i.e., $T(s,a,s^{\prime})=P(s^{\prime}|s,a)$.
POMDP [4, 41]: The agent maintains a belief over the partially-observable state $s$. It must infer the state using the observations $o$. Conversely, in a GHP-MDP, we do not have uncertainty over the observable part of the state $s$, and the uncertainty arises from the unknown latent factors $\theta$ that are affecting the environment dynamics.
Hidden Model MDP, Contextual MDP (cMDP) or Latent-Variable MDP (LMDP) [13, 36, 54, 87]: The model consists of a set of transition and reward functions from the same state and action space. The number of transition and reward functions is given by the number of contexts the agent could be in and the value function is the expectation over these, weighted according to a set of mixing weights. On the other hand, GHP-MDPs have a single transition and reward function whose input is augmented by the latent representation.
@@ III-B Model-Based Reinforcement Learning
An MBRL agent interacts with the environment by learning its model and using it to make decisions. A key difference with other system identification techniques is the exploratory nature of MBRL. Specifically, the agent acts on the environment while balancing exploration and exploitation [89]. Every episode it visits more informative regions as it avoids what it expects to be known or irrelevant configurations of the state-action space, which translates to faster model learning via informed data collection. This approach contrasts with MFRL, where the agent directly learns a policy solely based on interactions without explicit knowledge of the environment’s dynamics.
Formally, an MBRL agent learns a model of the environment, defined by a transition function $\hat{T}(s^{\prime}|s,a)$ and a reward function $\hat{R}(s,a)$, which approximate the true environment dynamics $T(s^{\prime}|s,a)$ and reward $R(s,a)$, respectively. To learn $\hat{T}$, the transition information =$\mathcal{R}=\{(s_{i},a_{i},s_{i+1})\}_{i=1}^{N}$ obtained through interaction and/or previously collected data is used for fine-tunning at the beginning of each episode. Specifically, the objective is to minimize the error between the predicted dynamics and the true dynamics of the environment:
@@
Then, during the episode the agent can simulate trajectories using the learned models $(\hat{T},\hat{R})$, and optimize the policy $\pi$ using any standard planning algorithm, such as value iteration or model predictive control (MPC). For example, to get the optimal action sequence from state $s_{t}$:
@@ III-C Bayesian Online Changepoint Detection (BOCD)
The problem of changepoint detection is concerned with determining the point where the observed data distribution changes in an ordered set of measurements, such as in time-series [25]. BOCD was introduced by Adams and MacKay [3] to tackle this problem by framing it as an estimation of the posterior distribution of the current “run length” $r_{t}$, meaning how likely it is that the measurement at time-step $t$ belongs to the same data generating process that started $r_{t}$ timesteps ago; while also obtaining the parameters $\boldsymbol{\eta}$ defining such process.
@@
where $H(\tau)$ is the hazard function [27]. Subsequently, finding the distribution of the run length $P(r_{t}|\boldsymbol{x}_{1:t})$, involves recursive computation over the run length and the observations:
@@ IV-A1 Multivariate BOCD:
Since we want to understand when the parameters $\boldsymbol{\eta}$ of the UDGP have changed, we need to define a prior for them. For this, we propose the Normal-Wishart distribution. This conjugate prior is particularly useful because it allows for closed form belief updates for $\boldsymbol{\mu}\in\mathbb{R}^{d}$ and $\Lambda\in\mathbb{R}^{d\times d}$ as more data becomes available. It is defined as follows [67]:
@@
Essentially, the Normal-Wishart prior combines two distributions: it leverages the Normal distribution for the mean $\boldsymbol{\mu}$, incorporating prior knowledge $\boldsymbol{\mu}_{0}$ about the likely location of $\boldsymbol{\mu}$, and it employs the Wishart distribution to model the precision matrix $\Lambda$, acting as the prior on the inverse covariance. The Wishart distribution is flexible, encoding varying levels of confidence in the scale of the precision matrix through its degrees of freedom parameter and scale matrix, $T$. Parameters $\kappa_{0}$ and $\nu_{0}$ allow us to control the strength of our prior beliefs, with $\kappa_{0}$ scaling the covariance of the mean and $\nu_{0}$ defining the degrees of freedom of the Wishart distribution. This prior choice is satisfactory for our application since it assumes an unknown mean vector $\boldsymbol{\mu}$ and precision matrix $\Lambda$, which enables flexibility by allowing uncertainty in the covariance structure to be encoded in a principled way [20, 67]. Consequently, it spares us from imposing restrictive assumptions on the covariance matrix, such as isotropy, homogeneity, or mean-dependence, making it highly suitable for modeling multivariate normal distributions with little assumptions about its parameters.
@@
Specifically, we will determine the closed form solutions for the growth probability Eq. (5) and changepoint probability Eq. (6), from which we can recover the run length probability Eq. (7). We begin by assuming that the prior on the changepoint probability is $P_{\mathtt{gap}}(g)\sim\mathtt{geometric}(\lambda)$. This makes the process memoryless and the hazard function constant at $H(\tau)=1/\lambda$, where $\lambda$ can be interpreted as the expected time between changepoints [3].
@@
    $P(r_{1}=1,$ $\leavevmode\nobreak\ \boldsymbol{x}_{0:1}=\{\boldsymbol{x}_{1}\})$
@@
The proof follows from the results of the posterior distribution parameters for the Normal-Wishart prior in [20, Sec. 9.10]. Specifically, they show that for a set of samples $\{\boldsymbol{x}_{i}:\boldsymbol{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{n}$, the posterior parameters for Eq. (11), given prior parameters $\{\boldsymbol{\mu}_{0},\kappa_{0},T_{0},\nu_{0}\}$, are:
@@
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \cdot\frac{P(r_{t}=\tau,\boldsymbol{x}_{0:t})}{\sum_{\begin{subarray}{c}\tau^{\prime}=0\end{subarray}}^{r_{t}}P(r_{t}=\tau^{\prime},\boldsymbol{x}_{0:t})},$ (18)
@@
From the BOCD framework, we assumed it is possible to compute the predictive distribution conditional on a given run length [3]:
@@
Additionally, $P(\boldsymbol{x}_{t+1}|\boldsymbol{x}_{1:t})$ depends only on the data of the current run $\boldsymbol{x}_{t}^{(r)}$. Thus, $P(\boldsymbol{x}_{t}|r_{t-1},\boldsymbol{x}_{1:t})=P(\boldsymbol{x}_{t}|r_{t-1},\boldsymbol{x}_{t}^{(r)})$. Therefore, using the result from [67] that says that the posterior predictive of the Normal-Wishart prior is:
@@ IV-B Online Situation Identification
To identify which situation $\psi\in\boldsymbol{\Psi}$ the robot is currently experiencing we employ the average Mahalanobis distance [64]:
@@
3:$\mathtt{situation\_change},\boldsymbol{\eta}_{t}\leftarrow\mathtt{SM}(\boldsymbol{x}_{t},\boldsymbol{x}^{(r)}_{t},\boldsymbol{\eta}_{t-1})$triangleright Alg. 2
4:$\psi^{*}\leftarrow\operatorname*{arg\,min}_{\psi\in\boldsymbol{\Psi}}M(\psi)$triangleright $M(\psi)=$ Eq. (19)
@@
7:         Add $(\boldsymbol{\mu},\Lambda)\leftarrow\text{MLE}(\boldsymbol{x}_{t}^{(r)})$ to $\boldsymbol{\Psi}$triangleright Create Situation
@@
10:    $\boldsymbol{x}_{t}^{(r)}\leftarrow\emptyset$triangleright Variable-sized window of past states
@@ IV-C Situationally-Aware Dynamics
In practical terms, we model the right-hand side of Eq. (20) as an ensemble $\mathcal{M}_{m}:S\times\Theta\times A\rightarrow S$ of $m$ probabilistic (Gaussian) neural networks. The ensemble’s input is augmented with a representation of the current situation (local transition dynamics $T_{L}$) of the robot, in order to inform it about the hidden state of the world. This architecture choice estimates the predictive distribution over the next state and simultaneously considers the epistemic and aleatoric uncertainty [15].
@@ IV-C1 Representing the Local Transition Model:
We find that the Moment Generating Function (MGF) provides a useful tool for this purpose, as it uniquely characterizes a distribution [18, Theorem 1]. In general, for a multivariate Gaussian random vector $\boldsymbol{X}\sim\mathcal{N}(\boldsymbol{\mu},\Sigma)$, the MGF is defined as
@@ IV-C2 Dynamics Learning:
The model $\mathcal{M}_{m}$ is learned by alternating exploration and exploitation, as is common in the MBRL framework. During the first episode, the agent’s actions are considered random as it is relying on a model initialized with random weights that has not yet learned anything about the environment. This exploration phase collects transition information for the replay buffer =$\mathcal{R}=\{(s_{i},\theta_{i},a_{i},s_{i+1})\}_{i=1}^{N_{e}}$, where $N_{e}$ is the number of steps in the episode. Note that different from traditional MBRL, we include the representation $\theta_{i}$ of the current situation of the robot in the replay buffer. The collected data is then used to train the dynamics model $\mathcal{M}_{m}$. In subsequent episodes, the agent uses the updated model to make better informed decisions, balancing exploration and exploitation. After each episode, the newly collected data is added to $\mathcal{R}$, and the model is retrained to refine its understanding of the environment’s dynamics. Throughout the training phase, the situation identification process is running in parallel to detect and model the new situations $\psi$, as detailed in Algorithm 3.
@@ IV-C3 Planning and Control:
We use our situationally-aware dynamics model $\mathcal{M}_{m}$ together with the Model Predictive Path Integral (MPPI) method [29]. The planning process involves generating a set of candidate action sequences over a time horizon $H$ and selecting the sequence that maximizes the expected return. The expected return is computed based on the predicted future states while assuming that the situation symbol $\theta$ remains constant throughout the rollouts. To select an action, MPPI optimizes the function:

### v2

@@ II-E Unstructured Terrain Navigation
The validation of our approach is through the challenging task of unstructured terrain navigation, where unmodeled and unmeasured terrain characteristics can significantly impact the robot’s motion. Unstructured terrain navigation has been tackled mostly from two perspectives: adaptation and path planning. Works addressing this from the adaptive perspective typically devise methods that allow the robot to recover from unexpected or hard scenarios. For example, xu2024kernel proposes a method to bypass the precise modeling of the state transition function of an MDP with continuous state space and shows that it can be applied for off-road navigation and recovering from unexpected phenomena (e.g., wheel slippage). wang2021rough learns a probabilistic dynamics model trained to consider the propagating effects of uncertainty and uses it along a tracking controller and constrained optimization to find trajectories with low divergence. Terrain-aware methods have been recently developed to generate consistent navigation behaviors by learning a control offset from demonstrations [siva2019robot, siva2021enhancing]. On the other hand, the methods concerned with global planning typically use the robot’s sensors to analyze the terrain and try to find the trajectory that seems safer and easier to navigate [yoo2024traversability, jardali2024autonomous, leininger2024gaussian]. Other methods have done so by focusing on risk-awareness [sharma2023ramp, cai2022risk, cai2023probabilistic]. Learning-based approaches have considered ground-robot interaction and surface data [sathyamoorthy2022terrapn]. Our method can be framed at the intersection of both perspectives. In particular, we obtain dynamics models with greater robustness to terrain challenges and use those to make local motion plans that help the robot traverse safely under partial state information.
@@ III-A Generalized Hidden Parameter Markov Decision Process (GHP-MDP)
Markov Decision Processes (MDPs) provide a framework for modeling decision-making with uncertainty over the action outcomes due to an agent’s interaction with its environment [sutton2018reinforcement]. Hidden Parameter MDPs (HiP-MDP) extend MDPs to scenarios where predicting the future also requires a set of unobservable parameters $\Theta$ that correspond to the sufficient statistics to adapt to a new situation [doshi2016hidden]. In HiP-MDPs, $\Theta$ explicitly represent a critical variable of the environment; for example, physical properties like the terrain friction. This makes it necessary to know a-priori which missing variables need to be estimated by $\Theta$, which is a strong assumption we cannot fulfill when the robot is navigating unstructured terrains as we cannot anticipate every challenge it will face.
[perez2020generalized] build upon HiP-MDPs and present GHP-MDPs to account for unobserved, latent parameters which influence the environment’s dynamics and reward function. These latent factors are not directly observable by the agent, but the agent must infer them through interaction with the environment to optimize its policy.
@@ III-A1 Comparison to other MDP Models:
MDP [bellman1958dynamic]: The state $s$ is fully observable, and the environment’s dynamics are completely determined by the state and action, i.e., $T(s,a,s^{\prime})=P(s^{\prime}|s,a)$.
POMDP [aastrom1965optimal, kaelbling1998planning]: The agent maintains a belief over the partially-observable state $s$. It must infer the state using the observations $o$. Conversely, in a GHP-MDP, we do not have uncertainty over the observable part of the state $s$, and the uncertainty arises from the unknown latent factors $\theta$ that are affecting the environment dynamics.
Hidden Model MDP, Contextual MDP (cMDP) or Latent-Variable MDP (LMDP) [chades2012momdps, hallak2015contextual, kwon2021rl, steimle2021multi]: The model consists of a set of transition and reward functions from the same state and action space. The number of transition and reward functions is given by the number of contexts the agent could be in and the value function is the expectation over these, weighted according to a set of mixing weights. On the other hand, GHP-MDPs have a single transition and reward function whose input is augmented by the latent representation.
@@ III-B Model-Based Reinforcement Learning
An MBRL agent interacts with the environment by learning its model and using it to make decisions. A key difference with other system identification techniques is the exploratory nature of MBRL. Specifically, the agent acts on the environment while balancing exploration and exploitation [sutton2018reinforcement]. Every episode it visits more informative regions as it avoids what it expects to be known or irrelevant configurations of the state-action space, which translates to faster model learning via informed data collection. This approach contrasts with MFRL, where the agent directly learns a policy solely based on interactions without explicit knowledge of the environment’s dynamics.
Formally, an MBRL agent learns a model of the environment, defined by a transition function $\hat{T}(s^{\prime}|s,a)$ and a reward function $\hat{R}(s,a)$, which approximate the true environment dynamics $T(s^{\prime}|s,a)$ and reward $R(s,a)$, respectively. To learn $\hat{T}$, the transition information $\mathcal{R}=\{(s_{i},a_{i},s_{i+1})\}_{i=1}^{N}$ obtained through interaction and/or previously collected data is used for fine-tunning at the beginning of each episode. Specifically, the objective is to minimize the error between the predicted dynamics and the true dynamics of the environment:
@@
[…]
@@ III-C Bayesian Online Changepoint Detection (BOCD)
The problem of changepoint detection is concerned with determining the point where the observed data distribution changes in an ordered set of measurements, such as in time-series [fearnhead2019changepoint]. BOCD was introduced by adams2007bayesian to tackle this problem by framing it as an estimation of the posterior distribution of the current “run length” $r_{t}$, meaning how likely it is that the measurement at time-step $t$ belongs to the same data generating process that started $r_{t}$ timesteps ago; while also obtaining the parameters $\boldsymbol{\eta}$ defining such process.
@@
where $H(\tau)$ is the hazard function [forbes2011statistical]. Subsequently, finding the distribution of the run length $P(r_{t}|\boldsymbol{x}_{1:t})$, involves recursive computation over the run length and the observations:
@@ IV-A1 Multivariate BOCD:
Since we want to understand when the parameters $\boldsymbol{\eta}$ of the UDGP have changed, we need to define a prior for them. For this, we propose the Normal-Wishart distribution. This conjugate prior is particularly useful because it allows for closed form belief updates for $\boldsymbol{\mu}\in\mathbb{R}^{d}$ and $\Lambda\in\mathbb{R}^{d\times d}$ as more data becomes available. It is defined as follows [murphy2007conjugateforms]:
@@
Essentially, the Normal-Wishart prior combines two distributions: it leverages the Normal distribution for the mean $\boldsymbol{\mu}$, incorporating prior knowledge $\boldsymbol{\mu}_{0}$ about the likely location of $\boldsymbol{\mu}$, and it employs the Wishart distribution to model the precision matrix $\Lambda$, acting as the prior on the inverse covariance. The Wishart distribution is flexible, encoding varying levels of confidence in the scale of the precision matrix through its degrees of freedom parameter and scale matrix, $T$. Parameters $\kappa_{0}$ and $\nu_{0}$ allow us to control the strength of our prior beliefs, with $\kappa_{0}$ scaling the covariance of the mean and $\nu_{0}$ defining the degrees of freedom of the Wishart distribution. This prior choice is satisfactory for our application since it assumes an unknown mean vector $\boldsymbol{\mu}$ and precision matrix $\Lambda$, which enables flexibility by allowing uncertainty in the covariance structure to be encoded in a principled way [degroot1970optimal, murphy2007conjugateforms]. Consequently, it spares us from imposing restrictive assumptions on the covariance matrix, such as isotropy, homogeneity, or mean-dependence, making it highly suitable for modeling multivariate normal distributions with little assumptions about its parameters.
@@
Specifically, we will determine the closed form solutions for the growth probability Eq. (5) and changepoint probability Eq. (6), from which we can recover the run length probability Eq. (7). We begin by assuming that the prior on the changepoint probability is $P_{\mathtt{gap}}(g)\sim\mathtt{geometric}(\lambda)$. This makes the process memoryless and the hazard function constant at $H(\tau)=1/\lambda$, where $\lambda$ can be interpreted as the expected time between changepoints [adams2007bayesian].
@@
    $P(r_{1}=1,$ $\penalty 10000\ \boldsymbol{x}_{0:1}=\{\boldsymbol{x}_{1}\})$
@@
The proof follows from the results of the posterior distribution parameters for the Normal-Wishart prior in [degroot1970optimal, Sec. 9.10]. Specifically, they show that for a set of samples $\{\boldsymbol{x}_{i}:\boldsymbol{x}_{i}\in\mathbb{R}^{d}\}_{i=1}^{n}$, the posterior parameters for Eq. (11), given prior parameters $\{\boldsymbol{\mu}_{0},\kappa_{0},T_{0},\nu_{0}\}$, are:
@@
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \cdot\frac{P(r_{t}=\tau,\boldsymbol{x}_{0:t})}{\sum_{\begin{subarray}{c}\tau^{\prime}=0\end{subarray}}^{r_{t}}P(r_{t}=\tau^{\prime},\boldsymbol{x}_{0:t})},$ (18)
@@
From the BOCD framework, we assumed it is possible to compute the predictive distribution conditional on a given run length [adams2007bayesian]:
@@
Additionally, $P(\boldsymbol{x}_{t+1}|\boldsymbol{x}_{1:t})$ depends only on the data of the current run $\boldsymbol{x}_{t}^{(r)}$. Thus, $P(\boldsymbol{x}_{t}|r_{t-1},\boldsymbol{x}_{1:t})=P(\boldsymbol{x}_{t}|r_{t-1},\boldsymbol{x}_{t}^{(r)})$. Therefore, using the result from [murphy2007conjugateforms] that says that the posterior predictive of the Normal-Wishart prior is:
@@ IV-B Online Situation Identification
To identify which situation $\psi\in\boldsymbol{\Psi}$ the robot is currently experiencing we employ the average Mahalanobis distance [mahalanobis1930test]:
@@
[…]
@@
7:    Add $(\boldsymbol{\mu},\Lambda)\leftarrow\text{MLE}(\boldsymbol{x}_{t}^{(r)})$ to $\boldsymbol{\Psi}$ $\triangleright$ Create Situation
@@
10:  $\boldsymbol{x}_{t}^{(r)}\leftarrow\emptyset$ $\triangleright$ Variable-sized window of past states
@@ IV-C Situationally-Aware Dynamics
In practical terms, we model the right-hand side of Eq. (20) as an ensemble $\mathcal{M}_{m}:S\times\Theta\times A\rightarrow S$ of $m$ probabilistic (Gaussian) neural networks. The ensemble’s input is augmented with a representation of the current situation (local transition dynamics $T_{L}$) of the robot, in order to inform it about the hidden state of the world. This architecture choice estimates the predictive distribution over the next state and simultaneously considers the epistemic and aleatoric uncertainty [chua2018deep].
@@ IV-C1 Representing the Local Transition Model:
We find that the Moment Generating Function (MGF) provides a useful tool for this purpose, as it uniquely characterizes a distribution [curtiss1942note, Theorem 1]. In general, for a multivariate Gaussian random vector $\boldsymbol{X}\sim\mathcal{N}(\boldsymbol{\mu},\Sigma)$, the MGF is defined as
@@ IV-C2 Dynamics Learning:
The model $\mathcal{M}_{m}$ is learned by alternating exploration and exploitation, as is common in the MBRL framework. During the first episode, the agent’s actions are considered random as it is relying on a model initialized with random weights that has not yet learned anything about the environment. This exploration phase collects transition information for the replay buffer $\mathcal{R}=\{(s_{i},\theta_{i},a_{i},s_{i+1})\}_{i=1}^{N_{e}}$, where $N_{e}$ is the number of steps in the episode. Note that different from traditional MBRL, we include the representation $\theta_{i}$ of the current situation of the robot in the replay buffer. The collected data is then used to train the dynamics model $\mathcal{M}_{m}$. In subsequent episodes, the agent uses the updated model to make better informed decisions, balancing exploration and exploitation. After each episode, the newly collected data is added to $\mathcal{R}$, and the model is retrained to refine its understanding of the environment’s dynamics. Throughout the training phase, the situation identification process is running in parallel to detect and model the new situations $\psi$, as detailed in Algorithm 3.
@@ IV-C3 Planning and Control:
We use our situationally-aware dynamics model $\mathcal{M}_{m}$ together with the Model Predictive Path Integral (MPPI) method [gandhi2021robust]. The planning process involves generating a set of candidate action sequences over a time horizon $H$ and selecting the sequence that maximizes the expected return. The expected return is computed based on the predicted future states while assuming that the situation symbol $\theta$ remains constant throughout the rollouts. To select an action, MPPI optimizes the function: